# On Partition Functions for Time-Inhomogeneous Branching Random Walks

### 15 Mar 2026 v1

@@ 3.1 Behavior of Partition Function at Criticality $\bar{\beta}_{c}$
Theorem 3.7.
(Asymptotics of partition function at criticality) At the critical parameter $\beta_{c}$, we have $\bar{W}_{n}^{\beta_{c}}=n^{-\frac{1}{2}+o(1)}$ almost surely.
@@
Our main goal in this step is see that it is suffice to show that the same asymptotics holds for all fractional moments $\alpha\in(0,1)$. The arguments here are fairly standard, and variations of these arguments are also employed in related works [12] and [2]:
@@
We abbreviate the notation by defining the ”partial free energy” for the time-inhomogeneous weight as:
@@
$\mathbb{P}\left(\bar{W}_{n}^{\beta_{c}}>e^{a_{n}-\epsilon|a_{n}|}\right)$ $\geq\Big(1-\frac{e^{\alpha(a_{n}-\epsilon|a_{n}|)}}{\mathbb{E}[(\bar{W}_{n}^{\beta_{c}})^{\alpha}]}\Big)^{2}\frac{\mathbb{E}[(\bar{W}_{n}^{\beta_{c}})^{\alpha}]^{2}}{\mathbb{E}[(\bar{W}_{n}^{\beta_{c}})^{2\alpha}]}$
$\geq\big(1-e^{-3\epsilon\alpha|a_{n}|/4}\big)^{2}e^{-\frac{3\epsilon}{4}|a_{n}|}\geq e^{-\epsilon|a_{n}|}\,$, (14)
@@
so that $\tau_{n}<n$ for all sufficiently large $n$ and $\tau_{n}\ll n$. We then have the decomposition:
$\bar{W}_{n}^{\beta_{c}}$ $=\sum_{|w|=\tau_{n}}\frac{1}{d^{\tau_{n}}}e^{\beta_{c}\bar{H}_{\tau_{n}}(w)-\frac{1}{2}\beta_{c}^{2}\sum_{i=1}^{\tau_{n}}f(\frac{i}{n})}\!\!\!\!\sum_{\begin{subarray}{c}v\in D_{n-\tau_{n}}(w)\end{subarray}}\frac{1}{d^{n-\tau_{n}}}e^{\beta_{c}\bar{H}_{\tau_{n}\to n}(w)-\frac{1}{2}\beta_{c}^{2}\sum_{i=\tau_{n}+1}^{n}f(\frac{i}{n})}$
@@
$\mathbb{P}\Big(\bar{W}_{n}^{\beta_{c}}\leq e^{a_{n}-\epsilon|a_{n}|}\exp\{\max_{|v|=\tau_{n}}\bar{V}_{\tau_{n}}^{n}(v)\}\Big)$ $\leq\mathbb{P}\Big(\sum_{|w|=\tau_{n}}Y_{n-\tau_{n}}(v)\leq e^{a_{n}-\epsilon|a_{n}|}\Big)$
$\leq\mathbb{P}\left(Y_{n-\tau_{n}}(v)\leq e^{a_{n}-\epsilon|a_{n}|}\right)^{d^{\tau_{n}}}$
Equation (14) implies that this expression is bounded above by $\exp\{-e^{\epsilon|a_{n}|}\}$ (since the fluctuations of the increments are decreasing). Therefore, by the assumption that $|a_{n}|\gg(\log n)^{\frac{1}{2}}$, these probabilities are summable. By the Borel-Cantelli lemma, it follows that almost surely
@@
Let $F$ be a concave function such that:
$\forall x\in\mathbb{R}_{+},|F(x)|\leq M\left(1+|x|^{a}\right)$
for some positive constants $M,a$, and let $\sigma$ be a Radon measure on the Borel subsets of $\mathbb{R}^{d}$. Given a bounded Borel set $A$, let $(X_{r})_{r\in A},(Y_{r})_{r\in A}$ be two continuous centered Gaussian processes with continuous covariance kernels $K_{X}$ and $K_{Y}$ such that:
@@
In this case, we may couple the law of the G-W tree for both time-homogeneous and time-inhomogeneous BRWs such that they have the same offspring distribution, and embed each realization of the G-W tree into the unit interval (i.e., $A=[0,1]$; see Section 2 for the detailed construction of such an embedding), and let the Borel measure $\sigma$ be the point mass associated to each realization. Note that for any $\alpha<1$, $F(x)=x^{\alpha}$ satisfies the conditions in the lemma.
@@
The upper bound of $\bar{W}_{n}^{\beta_{c}}$ follows by a similar argument as in Proposition 3.6, except that Hölder’s inequality is used here instead of the Cauchy-Schwarz inequality. We need to show that for any $\epsilon^{\prime}>0$, we have $\mathbb{E}[(\bar{W}_{n}^{\beta})^{\alpha}]\leq n^{-\frac{1}{2}+\epsilon^{\prime}}$. For all $p,q>1$ satisfying $1/p+1/q=1$, and recalling the expectation $\mathbf{E}_{n_{1}}$ under the (random) measure $\bar{\mathbb{P}}^{n_{1}}$ defined in Section 11, we have:
@@
$\mathbb{E}\left[(\bar{W}_{n}^{\beta})^{\alpha}\right]\leq\mathbb{E}\left[(\bar{W}_{n}^{n^{1-\epsilon}})^{p\alpha}\right]^{1/p}\mathbb{E}\left[\left(\mathbf{E}_{n^{1-\epsilon}}\left[\tilde{\bar{W}}_{n}^{n^{1-\epsilon}}\right]\right)^{q\alpha}\right]^{1/q}$
@@
Note that for any $q>1$, the second term on the right-hand side of the above inequality equals 1. As a result, for any fixed $\alpha>0$, $n_{1}>0$ and $\epsilon>0$, we may let $p$ be arbitrarily close to 1 so that $p\alpha<1$. Since $f$ is continuous at 0, for any fixed $\epsilon>0$, when $n$ is large enough, there exists $\delta>0$ such that $\sum_{i=1}^{n^{1-\epsilon}}\left(f(\frac{i}{n})-1\right)<(1-\delta)n$. By Kahane’s concentration inequality (Lemma 3.10), we can see that:
@@
$\mathbb{E}[(W_{(1-\delta)n}^{\beta})^{p\alpha}]\geq\mathbb{E}\left[(\bar{W}_{n}^{n^{1-\epsilon}})^{p\alpha}\right]$
@@
$\mathbb{E}\left[(\bar{W}_{n}^{n^{1-\epsilon}})^{p\alpha}\right]\leq((1-\delta)n)^{-\frac{1}{2p}+o(1)}\leq n^{-\frac{1}{2p}(1-\delta)+o(1)}$
@@
$\limsup_{n\to\infty}\mathbb{E}\left[(\bar{W}_{n}^{\beta})^{\alpha}\right]\leq n^{-\frac{1}{2p}(1-\delta)+o(1)}$
@@ 4 Further Directions
4.1 Behavior of Partition Function at Criticality $\bar{\beta}_{c}$
The derivative martingale is another key martingale associated with branching random walks and plays a central role in the study of their extreme values (see Section 5.4 of [17] for more details). At the critical parameter $\beta=\beta_{c}$, Biggins and Kyprianou [5] showed that the derivative martingale (as the name suggests, by taking ”derivative” with respect to $\beta$ of $W_{n}^{\beta}$ at $\beta_{c}$), defined by:
@@ 4.2 Relationship with Two-Dimensional Directed Polymers
There is an analogy between two-dimensional directed polymers and branching random walks with decreasing variance, and the former serves as one of the primary motivation for the later model discussed in the current paper.
Directed polymers in random environment describe the behavior of a long directed chain of monomers in presence of random impurities, see [8] for a comprehensive introduction to this area. In this setting, the trajectory of the polymer is given by a nearest-neighbor path $(S_{n})_{n\in\mathbb{N}}$ on two-dimensional lattice, while the impurities (also called the environment) are given by i.i.d. centered random variables $\omega(n,x)$, $n\in\mathbb{N},x\in\mathbb{Z}^{d}$ of variance one satisfying the free energy $\lambda(\beta)=\log\mathbb{E}[e^{\beta\omega(n,x)}]\in(0,\infty)$ for all $\beta>0$. The polymer measure on the path space is then defined by:
$\text{d}P_{N,\beta}^{\omega}(S)=\frac{e^{\beta\sum_{n=1}^{N}\left(\omega(n,S_{n})-\lambda(\beta)\right)}}{Z_{N}(\beta)}\text{d}P(S),$ (16)
where $\beta>0$ is the inverse temperature, $P=P_{0}$ with $(P_{x},E_{x})$ the probability measure and expectation of the simple random walk on $\mathbb{Z}^{d}$ started at point $x\in\mathbb{Z}^{d}$, and $Z_{N}(\beta)=E[e^{\beta\sum_{n=1}^{N}\left(\omega(n,S_{n})-\lambda(\beta)\right)}]$ is the (normalized) partition function.
For $x\in\mathbb{Z}^{2}$, define $W_{N}(x,\beta)=\mathbb{E}_{x}\left[e^{\sum_{n=1}^{N}\left(\beta\omega(n,S_{n})-\lambda(\beta)\right)}\right]$ to be the partition function for paths start at $x$. In the siminal paper [6], Caravenna, Sun and Zygouras established the convergence of the spatial behavior of two-dimensional partition functions:
Theorem 4.2.
([6], Theorem 2.5) Assume $d=2$ and set $\beta_{N}$ defined above. For all $\hat{\beta}\geq 1$ one has $W_{N}(x,\beta_{N})$ converges to 0 in probability as $N\to\infty$, while the log-partition function satisfies a central limit theorem in the weak disorder regime ($\hat{\beta}<1$):
$\forall\hat{\beta}\in(0,1),\log W_{N}(x,\beta_{N})\to^{(d)}\mathcal{N}(-\frac{\lambda^{2}}{2},\lambda^{2}),\text{ where }\lambda^{2}:=\log(\frac{1}{1-\hat{\beta}^{2}})$
In [9], the authors showed that when $\omega(n,x)$ are i.i.d. standard Gaussian (i.e. having law $\mathcal{N}(0,1)$), where $\lambda(\beta)=\frac{1}{2}\beta^{2}$, the normalized partition function for two-dimensional directed polymers at $x\in\mathbb{Z}^{2}$:
$h_{N}(x):=\sqrt{\log N}\left(\log W_{N}(x\sqrt{N},\beta_{N})-\mathbb{E}[\log W_{N}(x\sqrt{N},\beta_{N})]\right)$
can be approximated by $M_{N}(x)=\sum_{k=1}^{M}X_{k}(x)$ for some $M>0$ (eventually large), where $X_{k}(x)$ is defined by the normalized ”$k$-th step” partition function:
$X_{k}(x):=\sqrt{\log N}\left(\log Z_{N}(x\sqrt{N},\beta_{N})-\mathbb{E}[\log Z_{N}(x\sqrt{N},\beta_{N})]\right)$
and the ”$k$-th step” partition function is defined by:
$Z_{N}(x,\beta_{N})=\mathbf{E}_{x}[e^{\sum_{n=t_{k-1}+1}^{t_{k}}\beta_{N}\left(w(n,S_{n})-\frac{1}{2}\beta^{2}\right)}],\text{ where $t_{k}=N^{\frac{k}{M}}$}$
Theorem 2.1 in [9] suggests that the limit (when $N\to\infty$) of increments $X_{k}$ are centered random variables with strictly increasing variance (in $k$). Heuristically, one can think of $M_{N}(x)$ as a branching random walk where the increments are given by the $X_{k}(x)$’s. Indeed, we can see that these increments are independent in $k$, satisfy that $X_{k}(x)$ and $X_{k}(y)$ are correlated (having high covariance) for small |xy||x-y| ($x,y\in\mathbb{Z}^{2}$), and decorrelate (having low variance) as |xy||x-y| grows large. In this analogy, the root of the tree corresponds to level $k=M$, and the leaves correspond to level $k=1$. Consequently, the asymptotic behavior of $h_{N}(x)$ should be governed by a (Gaussian) branching random walk model with decreasing variance.
It would be interesting to establish a rigorous connection between the time-inhomogeneous BRWs with decreasing variance studied in this paper and the log-partition function of two-dimensional directed polymers. In fact, this heuristic has been confirmed in [10] for the extreme values of $h_{N}(x)$, and raises an intriguing open problem: whether the two-dimensional polymer measure defined by:

### 27 Mar 2026 v2

@@ 3.3 Behavior of Partition Function at Criticality
Theorem 3.9 (Asymptotics of partition function at criticality).
At the critical parameter $\beta_{c}$, we have $\bar{W}_{n}^{\beta_{c}}=n^{-\frac{1}{2}+o(1)}$ almost surely.
@@
First, we will see that it is suffice to show the same asymptotics holds for all fractional moments for $\alpha\in(0,1)$. The arguments here are fairly standard, and variations of these arguments are also employed in related works [15] and [2]:
@@
We abbreviate the notation by defining the “partial free energy” for the time-inhomogeneous weight as (recall the definition of $\bar{H}_{n}^{\tau_{n}}(w)$ in the proof of Theorem 3.7):
@@
$\mathbb{P}\left\{\bar{W}_{n}^{\beta_{c}}>e^{a_{n}-\epsilon|a_{n}|}\right\}$ $\geq\left(1-\frac{e^{\alpha(a_{n}-\epsilon|a_{n}|)}}{\mathbb{E}[(\bar{W}_{n}^{\beta_{c}})^{\alpha}]}\right)^{2}\frac{\mathbb{E}[(\bar{W}_{n}^{\beta_{c}})^{\alpha}]^{2}}{\mathbb{E}[(\bar{W}_{n}^{\beta_{c}})^{2\alpha}]}$
$\geq\left(1-e^{-3\epsilon\alpha|a_{n}|/4}\right)^{2}e^{-\frac{3\epsilon}{4}|a_{n}|}\geq e^{-\epsilon|a_{n}|}\,$, (15)
@@
so that $\tau_{n}\ll n$ for all sufficiently large $n$. We then have the decomposition:
$\bar{W}_{n}^{\beta_{c}}$ $=\sum_{|w|=\tau_{n}}\frac{1}{d^{\tau_{n}}}e^{\beta_{c}\bar{H}_{n}^{\tau_{n}}(w)-\frac{1}{2}\beta_{c}^{2}\sum_{i=1}^{\tau_{n}}f(\frac{i}{n})}\!\!\!\!\sum_{\begin{subarray}{c}v\in D_{n-\tau_{n}}(w)\end{subarray}}\frac{1}{d^{n-\tau_{n}}}e^{\beta_{c}\bar{H}_{\tau_{n}\to n}(w)-\frac{1}{2}\beta_{c}^{2}\sum_{i=\tau_{n}+1}^{n}f(\frac{i}{n})}$
@@
$\mathbb{P}\left\{\bar{W}_{n}^{\beta_{c}}\leq e^{a_{n}-\epsilon|a_{n}|}\exp\{\max_{|v|=\tau_{n}}\bar{V}_{\tau_{n}}^{n}(v)\}\right\}$ $\leq\mathbb{P}\left\{\sum_{|w|=\tau_{n}}Y_{n-\tau_{n}}(v)\leq e^{a_{n}-\epsilon|a_{n}|}\right\}$
$\leq\mathbb{P}\left\{Y_{n-\tau_{n}}(v)\leq e^{a_{n}-\epsilon|a_{n}|}\right\}^{d^{\tau_{n}}}$
Equation (15) implies that this expression is bounded above by $\exp\{-e^{\epsilon|a_{n}|}\}$ (since the fluctuations of the increments are decreasing). Therefore, by the assumption that $|a_{n}|\gg(\log n)^{\frac{1}{2}}$, these probabilities are summable. By the Borel-Cantelli lemma, it follows that almost surely:
@@
(Kahane’s Concentration Inequality) Let $F$ be a concave function such that:
$\forall x\in\mathbb{R}_{+},|F(x)|\leq M\left(1+|x|^{m}\right)$
for some positive constants $M,m$, and let $\sigma$ be a Radon measure on the Borel subsets of $\mathbb{R}^{d}$. Given a bounded Borel set $A$, let $(X_{r})_{r\in A},(Y_{r})_{r\in A}$ be two continuous centered Gaussian processes with continuous covariance kernels $K_{X}$ and $K_{Y}$ such that:
@@
In this case, we may couple the law of the G-W tree for both time-homogeneous and time-inhomogeneous BRWs such that they have the same offspring distribution, and embed each realization of the G-W tree into the unit interval (i.e., $A=[0,1]$; see Section 2 for the detailed construction of such an embedding), and let the Borel measure $\sigma$ be the point mass associated to each realization. Note that for any $0<\alpha<1$, $F(x)=x^{\alpha}$ satisfies the conditions in the Lemma 3.12.
@@
The upper bound of $\bar{W}_{n}^{\beta_{c}}$ follows by a similar argument as in Proposition 3.7, except that Hölder’s inequality is used here instead of the Cauchy-Schwarz inequality. We need to show that for any $\bar{\epsilon}>0$, we have $\mathbb{E}[(\bar{W}_{n}^{\beta})^{\alpha}]\leq n^{-\frac{1}{2}+\bar{\epsilon}}$. For all $p,q>1$ satisfying $1/p+1/q=1$, and recalling the expectation $\mathbf{E}_{\lfloor n^{1-\epsilon}\rfloor}$ under the (random) measure $\bar{\mathbb{P}}^{\lfloor n^{1-\epsilon}\rfloor}$ defined in 12, we have:
@@
$\mathbb{E}\left[(\bar{W}_{n}^{\beta})^{\alpha}\right]\leq\mathbb{E}\left[(\bar{W}_{n}^{\lfloor n^{1-\epsilon}\rfloor})^{p\alpha}\right]^{1/p}\mathbb{E}\left[\left(\mathbf{E}_{\lfloor n^{1-\epsilon}\rfloor}\left[\tilde{\bar{W}}_{n}^{\lfloor n^{1-\epsilon}\rfloor}\right]\right)^{q\alpha}\right]^{1/q}$
@@
Note that for any $q>1$, the second term on the right-hand side of the above inequality equals 1. As a result, for any fixed $\alpha>0$ and $\epsilon>0$, we may let $p$ be arbitrarily close to 1 so that $p^{2}\alpha<1$. Since $f$ is continuous at 0, for any fixed $\epsilon>0$, when $n$ is large enough, there exists $\delta>0$ such that $\lfloor(1-\delta)n\rfloor<\sum_{i=1}^{\lfloor n^{1-\epsilon}\rfloor}\left(f(\frac{i}{n})-1\right)<n$. Then we can see that (where the inequality 17 follows from Kahane’s concentration inequality 3.12):
@@
$\mathbb{E}\left[(\bar{W}_{n}^{n^{1-\epsilon}})^{p\alpha}\right]=\mathbb{E}\left[\left(\frac{1}{d^{n}}\sum_{\mathbf{|x|}=n}e^{\beta_{c}\bar{H}_{n}^{\lfloor n^{1-\epsilon}\rfloor}(\mathbf{x})-\frac{1}{2}\sum_{i=1}^{\lfloor n^{1-\epsilon}\rfloor-1}\beta_{c}^{2}}\right)^{p\alpha}\right]$
$\leq\mathbb{E}\left[\left(\frac{1}{d^{n}}\sum_{|\mathbf{x}|=n}e^{(1-\delta)\beta_{c}H_{n}(\mathbf{x})-\frac{1}{2}\lfloor(1-\delta)n\rfloor\beta_{c}^{2}}\right)^{p\alpha}\right]$ (17)
$\leq\mathbb{E}\left[\left(\frac{1}{d^{\lfloor(1-\delta)n\rfloor}}\sum_{|\mathbf{x}|=\lfloor(1-\delta)n\rfloor}e^{\beta_{c}H_{\lfloor(1-\delta)n\rfloor}(\mathbf{x})-\frac{1}{2}\lfloor(1-\delta)n\rfloor\beta_{c}^{2}}\right)^{p^{2}\alpha}\right]^{1/p}\mathbb{E}\left[\left(\mathbf{E}_{\lfloor(1-\delta)n\rfloor}[\tilde{W}_{n}^{\lfloor(1-\delta)n\rfloor}]\right)^{pq\alpha}\right]^{1/q}$
$\mathbb{E}\left[\left(\bar{W}_{\lfloor(1-\delta)n\rfloor}^{\beta}\right)^{p^{2}\alpha}\right]^{1/p}\times 1$
where the inequality Kahane’s concentration inequality (Lemma 3.12),in the expressions above, analogous to that in the proof of Theorem 3.7, we define $\mathbf{E}_{\lfloor(1-\delta)n\rfloor}$ to be the expectation under the transformed measure $\bar{\mathbb{P}}^{\lfloor(1-\delta)n\rfloor}$ (recall $\mathbf{P}^{*}$ is the survival probability of GW tree):
$\frac{\mathrm{d}\bar{\mathbb{P}}^{\lfloor(1-\delta)n\rfloor}(\mathbf{x})}{\mathrm{d}\mathbf{P}^{*}}=\frac{e^{-\frac{1}{2}\beta^{2}\lfloor(1-\delta)n\rfloor+\beta H_{n}^{\lfloor(1-\delta)n\rfloor}(\mathbf{x})}}{W_{n}^{\lfloor(1-\delta)n\rfloor}}$ (18)
The “first step” partition function denoted $W_{n}^{\lfloor(1-\delta)n\rfloor}$ is defined as:
$\tilde{W}_{n}^{\lfloor(1-\delta)n\rfloor}=\frac{1}{d^{\lfloor(1-\delta)n\rfloor}}\sum_{|\mathbf{x}|=n}e^{\beta_{c}H_{n}^{\lfloor(1-\delta)n\rfloor}(\mathbf{x})-\frac{1}{2}\lfloor(1-\delta)n\rfloor\beta_{c}^{2}}\text{ where }H_{n}^{\lfloor(1-\delta)n\rfloor}(\mathbf{x})=\sum_{i=1}^{\lfloor(1-\delta)n\rfloor}w(x_{i})$
and its “complement” denoted $\tilde{\bar{W}}_{n}^{\lfloor(1-\delta)n\rfloor}$ is defined as:
$\tilde{\bar{W}}_{n}^{\lfloor(1-\delta)n\rfloor}=\frac{1}{d^{\lceil\delta n\rceil}}\sum_{\mathbf{y}\in D_{\lfloor(1-\delta)n\rfloor}^{\mathbf{x}}}e^{\beta_{c}H_{n}^{\lfloor(1-\delta)n\rfloor\to n}(\mathbf{y})-\frac{1}{2}\lceil\delta n\rceil\beta_{c}^{2}},\text{ where }H_{n}^{\lfloor(1-\delta)n\rfloor\to n}(\mathbf{x})=\sum_{i=\lfloor(1-\delta)n\rfloor}^{n}w(x_{i})$
@@
$\mathbb{E}\left[(\bar{W}_{n}^{n^{1-\epsilon}})^{p^{2}\alpha}\right]^{1/p}\leq\mathbb{E}\left[\left(\bar{W}_{\lfloor(1-\delta)n\rfloor}^{\beta}\right)^{p^{2}\alpha}\right]^{1/p^{2}}\leq\left(\lfloor(1-\delta)n\rfloor\right)^{-\frac{1}{2p^{2}}+o(1)}\leq n^{-\frac{1}{2p^{2}}(1-\delta)+o(1)}$
@@
$\limsup_{n\to\infty}\mathbb{E}\left[(\bar{W}_{n}^{\beta})^{\alpha}\right]\leq n^{-\frac{1}{2p^{2}}(1-\delta)+o(1)}$
@@ 4 Further Directions
4.1 Behavior of Partition Function at Criticality
The derivative martingale is another key martingale associated with branching random walks and plays a central role in the study of their extreme values (see Section 5.4 of [24] for more details). At the critical parameter $\beta=\beta_{c}$, Biggins and Kyprianou [5] showed that the derivative martingale (as the name suggests, by taking “derivative” with respect to $\beta$ of $W_{n}^{\beta}$ at $\beta_{c}$), defined by:
@@ 4.2 Relationship with Two-Dimensional Directed Polymers
Recall that we discussed the heuristic relationship between two-dimensional directed polymers and BRWs with decreasing variance in Section 2; it would be interesting to establish a rigorous connection between the time-inhomogeneous BRWs with decreasing variance studied in this paper and the log-partition function of two-dimensional directed polymers. In fact, this heuristic has been confirmed in [12] for the extreme values of $h_{N}(x)$ (the partition function under certain scaling and normalization), and raises an intriguing open problem: whether the two-dimensional polymer measure defined by
@@
4.3 GMC Beyond Log-correlation
An interesting implication of Theorem 3.3 is that if we embed the (binary) BRW into the unit interval $[0,1]$, the covariance structure $K(x,y)$ (for $x,y\in[0,1]$) differs from that of GMC (which takes the form $\log|x-y|+O(1)$): it is not log-correlated. The universality behavior we established enables us to investigate whether we can construct a GMC-type object beyond log-correlated fields, which defines a variant of GMC measure.